# A fully relativistic treatment of EMRIs in collisionless environments

### v1

@@ Abstract
Figure 2 characterizes the importance of relativistic effects in the modeling of environmental effects in terms of what future mHz interferometers like LISA can discern. The ability of the detector to distinguish two waveforms grows with the observation time as the SNR accumulates. In particular, one can estimate the critical mismatch above which two waveforms can be distinguished in the detector by $\mathcal{M}_{c}=(n/2)\,\mathrm{SNR}^{-2}$ [92], where $n$ is the number of intrinsic parameters (here, equal to $15$). The critical mismatch is shown in Fig. 2 as a function of the observation time for different luminosity distances (or optimal SNR at $2$ years of observation). For periods of observation with $\mathcal{M}>\mathcal{M}_{c}$, the relativistic effects must be included in the modeling of environmental effects to avoid systematic biases or risk a substantial loss in SNR. As a reference, for our benchmark system with $q=10^{-5}$, this observation period falls roughly between 1 and 4.5 months for luminosity distances $D_{L}\in(0.3,2.6)\,{\rm Gpc}$.
@@
Both the Newtonian and phenomenological models (with heuristic relativistic effects) can largely underestimate the impact of environmental effects on GW waveforms. For our benchmark system with mass-ratio $q=10^{-5}$, we find that a few years of observation before merger are required for LISA to discern those models from vacuum (see Supplemental Material), as opposed to the few weeks for the fully relativistic model. Our results also show that relativistic corrections to environmental effects are comparatively weaker for systems with less extreme mass ratios, which can be understood from the fact that such systems chirp faster through the strong-field region.
@@
Discussion.— In this letter, we introduce a fully relativistic framework to include environmental effects of collisionless environments in the evolution of EMRIs. Applying it to the case of DM spikes around massive BHs, we showed that a fully relativistic treatment is necessary to accurately describe the environmental dephasing on such signals. In this work, we neglected the effect of the EMRI on the DM overdensity. Using the Newtonian framework HaloFeedback [93, 27], we confirmed that an EMRI with mass-ratio $q=10^{-5}$ would lead to tiny feedback on the DM overdensity: the evolution of the distribution function would lead to less than one cycle of dephasing in 4 years (corresponding to $\sim 0.5\%$ of the total dephasing).
@@
A different route to relativistic environmental effects on the generation and propagation of GWs was taken in [34] (with subsequent works [35, 36, 37, 38, 39]), which includes the self-gravity of surrounding matter. This was based on a result by Einstein [94], that a system of many gravitating point masses moving along circular geodesics about a common center can be mapped into an effective anisotropic fluid with only tangential pressure. Because real DM halos should contain particles distributed over a wide range of eccentricities due to processes like violent relaxation [95, 96], the physical interpretation of results derived for “Einstein clusters” should be done with care. Very rich phenomenology is observed to emerge from the coupled (linear) perturbations in the gravitational-matter sectors in such approaches. Typical DM halos, as our benchmark MW-like halo, have dimensionless compactness values $\lesssim 10^{-6}$, leading to very small couplings between gravity and matter perturbations; it is conceivable, though, that such small effects might accumulate over the large halo scales into observable imprints in GW waveforms. In low-compactness collisionless halos, effects on GW propagation described within geometric optics approximation via a gravitational redshift [97] lead roughly to one dephasing cycle after $\sim 10^{6}$ cycles; this is $\sim 10^{3}$ times smaller than the effects of dynamical friction studied here (see also [55]).
@@
In this work, we imposed a hard cut-off on the impact parameter, neglecting (conservatively) all the interactions with DM particles outside the secondary’s Hill radius, where the validity of our assumptions breaks down. We expect the total torque resulting from additional gravitational three-body interactions to slightly increase the impact of environmental effects on EMRI evolution. Such “stirring” effect could, in principle, be obtained perturbatively by computing the energy and angular momentum exchange with the DM particles mediated by (vacuum) linear metric perturbations, similarly to the approach of [42] to ultralight scalar fields.
@@
Some direct extensions of this work include the application of our framework to eccentric and inclined EMRIs onto a primary Kerr BH. We note that cold DM spikes can be significantly more dense around spinning BHs [75]. Moreover, we expect relativistic effects in the interaction of EMRIs with environments to be even more important for prograde orbits around highly spinning BHs, as their ISCO (and marginally bound orbits) approaches the event horizon radius. Here, we neglected the spin of the secondary, but the results of [83] could be directly incorporated into our framework. Surrounding environments are responsible for a characteristic co-evolution of the mass and spin of the companion (to be explored in [98]); while unmeasurable for quasi-circular EMRIs [99], the spin of the companion might be probed with eccentricity [100].
Acknowledgements.
@@
We thank Francisco Duque and Lorenzo Speri for valuable discussions and guidance on the implementation of our results in FEW. We are also grateful to Miguel R. Correia for insightful conversations in the early stages of this work. We also thank Vitor Cardoso, Francisco Duque, and Andrea Maselli for their helpful comments and stimulating discussions on our draft. We gratefully acknowledge the support of the Dutch Research Council (NWO) through an Open Competition Domain Science-M grant, project number OCENW.M.21.375.
I Supplemental Material
Relativistic Coulomb logarithm
In computing the momentum transferred by particles scattering off a Schwarzschild spacetime, the contribution from orbits with asymptotic velocity $v$ and impact parameter $b\gtrsim b_{\rm cr}(v)$ (i.e., only slightly larger than the one of capture) must be computed numerically [32]. This contribution can be encapsulated in the factor $\chi(v)$ defined in the main text. Choosing $b_{\rm min}=40b_{\rm cr}(v)$, the function $\chi$ is accurately described (with an error $\leq 1.5\%$) by the fit
$\chi\approx\ln\sqrt{\tfrac{b_{\rm cr}^{2}(v)+b_{\rm gr}^{2}(v)}{b_{\rm min}^{2}+b_{\rm gr}^{2}(v)}}+12786v^{2}\left(\frac{1-0.096v^{3}}{1+1844v^{1.62}}\right)\,.$
DM distribution function in velocity space
To better understand the distribution of velocities in the companion’s frame, in Fig. 4 we show a density plot of the distribution function $\mathrm{d}N/(\mathrm{d}V\mathrm{d}^{3}v^{i})$ for a companion in a circular orbit [^6] Or at pericenter or apocenter. at the ISCO radius. We show the region $v^{1}\geq 0$ of a slice $v^{2}=0$, as the distribution function is symmetric under $v^{1}\to-v^{1}$ (for equatorial orbits, it is also symmetric under $v_{2}\to-v^{2}$).
The white region represents the loss cone of the DM spike, i.e., only particles with velocities within the colored regions are in stable bound orbits around the supermassive BH. All the others are either unbound or captured by the supermassive BH. We notice two disconnected regions associated with DM particles in prograde and retrograde motion relative to the companion orbit. While for a companion in circular orbit, at large radii the prograde region is roughly centered at $v^{3}=0$, we note that for orbital radii in the strong-field most prograde orbits have $v^{3}<0$, contributing a net positive torque to the EMRI. Simplified models neglecting the velocity anisotropy of the distribution cannot capture this radius-dependent competition between positive and negative torques, which can affect the final torque.
As explained in the text, both in the Newtonian and phenomenological (with heuristic relativistic effects) models, we used the distribution function perceived by stationary observers $\bm{n}=(-g_{tt})^{-\nicefrac{{1}}{{2}}}\bm{\partial}_{t}$ to evaluate the integral in $\mathrm{d}^{3}v^{i}$. Their different local plane of simultaneity changes the distribution function relatively to the companion’s (shown in Fig. 4 for $R=6M$) by a multiplicative factor $(-g_{tt})^{-\nicefrac{{1}}{{2}}}\mathcal{E}/\gamma_{v}$, while not changing the loss cone. This implies that these observers perceive a considerably smaller density of particles in retrograde orbits, and slightly larger in prograde orbits.
Figure 4: Distribution function in the companion frame at the ISCO ($R=6M$) for the region $v^{1}\geq 0$ of a slice $v^{2}=0$. For a (quasi)circular EMRI, the distribution function is symmetric under $v^{1}\to-v^{1}$. The white region shows the loss cone of our fiducial DM spike; only the velocities in the colored regions correspond to stable bound orbits around the supermassive BH.
Comparison between different models
In Fig. 5, we compare the faithfulness between multiple models as a function of observation period. The oscillations are due to the incidental interference for long waveforms and do not carry any physical meaning (see, e.g., [52]). We see that the Newtonian and phenomenological models (with heuristic relativistic effects) can largely underestimate the impact of environmental effects on GW waveforms. We can also notice that relativistic corrections to interaction with the environment are comparatively weaker for systems with less extreme mass ratios, because such systems chirp faster through the strong-field region.
Figure 5: Faithfulness between different models as a function of signal duration (in years before merger). The solid lines are for the benchmark $q=10^{-5}$ system, while the dashed are for a slightly heavier companion.

### v2

@@ Abstract
Figure 2 characterizes the importance of relativistic effects in the modeling of environmental effects in terms of what future mHz interferometers like LISA can discern. The ability of the detector to distinguish two waveforms grows with the observation time as the SNR accumulates. In particular, one can estimate the critical mismatch above which two waveforms can be distinguished in the detector by $\mathcal{M}_{c}=(n/2)\,\mathrm{SNR}^{-2}$ [99], where $n$ is the number of intrinsic parameters (here, equal to $15$). The critical mismatch is shown in Fig. 2 as a function of the observation time for different luminosity distances (or optimal SNR at $2$ yr of observation). For periods of observation with $\mathcal{M}>\mathcal{M}_{c}$, the relativistic effects must be included in the modeling of environmental effects to avoid systematic biases or risk a substantial loss in SNR. As a reference, for our benchmark system with $q=10^{-5}$, this observation period falls roughly between 1 and 4.5 months for luminosity distances $D_{L}\in(0.3,2.6)\,{\rm Gpc}$. While the precise values of the mismatch between the relativistic and Newtonian models may vary with the EMRI parameters, we do not expect our qualitative conclusions to be affected.
@@
Newtonian models largely underestimate the impact of environmental effects on EMRI waveforms. For our benchmark system with mass-ratio $q=10^{-5}$, we verified that an observation of more than a year before merger is required for LISA to discern those models from vacuum, as opposed to the few weeks for the fully relativistic model. Our results also show that relativistic corrections to environmental effects are comparatively weaker for systems with less extreme mass ratios, which can be understood from the fact that such systems chirp faster through the strong-field region.
@@
Discussion.— In this Letter, we introduce a fully relativistic framework to include environmental effects of collisionless environments in the evolution of EMRIs. Applying it to the case of DM spikes around massive BHs, we showed that a fully relativistic treatment is necessary to accurately describe the environmental dephasing on such signals. In this work, we neglected the effect of the EMRI on the DM overdensity. Using the Newtonian framework HaloFeedback [100, 29], we confirmed that an EMRI with mass-ratio $q=10^{-5}$ would lead to tiny feedback on the DM overdensity: the evolution of the distribution function would lead to less than one cycle of dephasing in 4 yr (corresponding to $\sim 0.5\%$ of the total dephasing). We note, though, that, for LISA data analysis, post-adiabatic EMRI waveforms with errors below one radian dephasing after 4 yr are required.
@@
A different route to relativistic environmental effects on the generation and propagation of GWs was taken in [36] (with subsequent works [37, 38, 39, 40, 41]), which includes the self-gravity of surrounding matter. This was based on a result by Einstein [101], that a system of many gravitating point masses moving along circular geodesics about a common center can be mapped into an effective anisotropic fluid with only tangential pressure. Because real DM halos should contain particles distributed over a wide range of eccentricities due to processes like violent relaxation [102, 103], the physical interpretation of results derived for “Einstein clusters” should be done with care. Very rich phenomenology is observed to emerge from the coupled (linear) perturbations in the gravitational-matter sectors in such approaches. Typical halos, as our benchmark Milky-Way-like halo, have dimensionless compactness $\lesssim 10^{-6}$, leading to very small couplings between gravity and matter perturbations; it is conceivable, though, that such small effects might accumulate over the large halo scales into observable imprints in GW waveforms. In low-compactness collisionless halos, effects on GW propagation described within geometric optics approximation via a gravitational redshift [104] lead roughly to one dephasing cycle after $\sim 10^{6}$ cycles; this is $\sim 10^{3}$ times smaller than the effects of dynamical friction studied here (as found in [57]).
@@
In this Letter, we imposed a hard cutoff on the impact parameter, (conservatively) neglecting the interactions with DM particles outside the secondary’s Hill radius, where the validity of our assumptions breaks down. We expect the total torque resulting from additional gravitational three-body interactions to increase the impact of environmental effects on EMRI evolution. Such “stirring” effect could, in principle, be obtained perturbatively by computing the energy and angular momentum exchange with the DM particles mediated by (vacuum) linear metric perturbations, similarly to the approach of [44] to ultralight scalar fields.
@@
Direct extensions of this work include the application of our framework to eccentric and inclined EMRIs onto a primary Kerr BH. We note that cold DM spikes can be significantly denser around spinning BHs [77]. Moreover, we expect relativistic effects in the interaction of EMRIs with environments to be even more important for prograde orbits around highly spinning BHs, as their ISCO (and marginally bound orbits) approaches the event horizon radius. Here, we neglected the spin of the secondary, but the results of [85] could be directly incorporated into our framework. Surrounding environments are responsible for a characteristic coevolution of the mass and spin of the companion (explored in [105]); while unmeasurable for quasicircular EMRIs [106], the spin of the companion might be probed with eccentricity [107].
@@
Acknowledgments.— We thank Francisco Duque and Lorenzo Speri for valuable discussions and guidance on implementing our results in FEW. We are also grateful to Miguel R. Correia for insightful conversations in the early stages of this work. We also thank Emanuele Berti, Vitor Cardoso, Francisco Duque, Andrea Maselli, and Nick Speeney for their helpful comments and stimulating discussions on our draft. We gratefully acknowledge the support of the Dutch Research Council (NWO) through an Open Competition Domain Science-M grant, project number OCENW.M.21.375.